0.001+0.02x-0.1=0.002x
Answer by BGluvsmath(64)

0.001+0.02x-0.1=0.002x
-0.02x -0.02x
----------------------
0.001-0.1 = -0.018x
-0.099 = -0.018x
5.5 = x (or x = 5.5)